Six new prototypes for internal-combustion engines
Answer #1
Concepts and reason

The concept used to solve this problem is work done by the heat engine.

First use the expression of the work done by the heat engine. Then compare the work done by all the given heat engine. Finally arrange the work done by heat engine from largest to smallest.

Fundamentals

The expression for the work done by the heat engine can be expressed as follows,

W=QinQoutW = {Q_{in}} - {Q_{out}} …… (1)

Here, Qin{Q_{in}} represents the heat energy taken from the source and Qout{Q_{out}} represents the heat energy taken from the system.

Calculate the work done by the heat engine 1.

Substitute 2000J2000{\rm{ J}} for Qin{Q_{in}} and 1000J1000{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W1=2000J1000J=1000J\begin{array}{c}\\{W_1} = 2000{\rm{ J}} - 1000{\rm{ J}}\\\\{\rm{ = 1000 J}}\\\end{array}

Calculate the work done by the heat engine 2.

Substitute 1500J{\rm{1500 J}} for Qin{Q_{in}} and 500J500{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W2=1500J500J=1000J\begin{array}{c}\\{W_2} = 1500{\rm{ J}} - 500{\rm{ J}}\\\\{\rm{ = 1000 J}}\\\end{array}

Calculate the work done by the heat engine 3.

Substitute 4000J{\rm{4000 J}} for Qin{Q_{in}} and 3000J3000{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W3=4000J3000J=1000J\begin{array}{c}\\{W_3} = 4000{\rm{ J}} - 3000{\rm{ J}}\\\\{\rm{ = 1000 J}}\\\end{array}

Calculate the work done by the heat engine 4.

Substitute 3000J{\rm{3000 J}} for Qin{Q_{in}} and 1500J1500{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W4=3000J1500J=1500J\begin{array}{c}\\{W_4} = 3000{\rm{ J}} - 1500{\rm{ J}}\\\\{\rm{ = 1500 J}}\\\end{array}

Calculate the work done by the heat engine 5.

Substitute 1000J{\rm{1000 J}} for Qin{Q_{in}} and 200J200{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W5=1000J200J=800J\begin{array}{c}\\{W_5} = 1000{\rm{ J}} - 200{\rm{ J}}\\\\{\rm{ = 800 J}}\\\end{array}

Calculate the work done by the heat engine 6.

Substitute 2000J{\rm{2000 J}} for Qin{Q_{in}} and 500J500{\rm{ J}} for Qout{Q_{out}} in the equation (1) W=QinQoutW = {Q_{in}} - {Q_{out}} .

W6=2000J500J=1500J\begin{array}{c}\\{W_6} = 2000{\rm{ J}} - 500{\rm{ J}}\\\\{\rm{ = 1500 J}}\\\end{array}

Compare the work done by all the heat engines.

W4,W6>W1,W2,W3>W5{W_4}{\rm{,}}{W_6}{\rm{ > }}{W_1},{W_2},{W_3}{\rm{ > }}{W_5}

Ans:

The rank of the work done by heat engines from largest to smallest is W4,W6>W1,W2,W3>W5{W_4}{\rm{,}}{W_6}{\rm{ > }}{W_1},{W_2},{W_3}{\rm{ > }}{W_5} .